OpenStudy (anonymous):

Please help! (2x^2-5x-3)/(x+1) What would be the slant asymptote for this?

OpenStudy (anonymous):

x=-1?

OpenStudy (anonymous):

How did you find that?

OpenStudy (nincompoop):

the limit should be able to tell you the asymptote

OpenStudy (anonymous):

How do you find the limit?

OpenStudy (nincompoop):

set x to approach -1 but not equal to -1

OpenStudy (nincompoop):

then tabulate the result

OpenStudy (anonymous):

So would it be x=-1?

OpenStudy (nincompoop):

what happens when you divide by zero?

OpenStudy (anonymous):

You can't divide by 0?

OpenStudy (nincompoop):

exactly... how about -.99?

OpenStudy (nincompoop):

or -0.9?

OpenStudy (anonymous):

To divide by?

OpenStudy (nincompoop):

-0.9 + 1

OpenStudy (anonymous):

So -1.9?

OpenStudy (nincompoop):

that would be the value you will use in place of x

OpenStudy (nincompoop):

you can go as close to -1 but not -1 itself

OpenStudy (nincompoop):

0.99999999999999999999999999999 etc

OpenStudy (nincompoop):

as you would go closer and closer to -1 you will see the value of asymptote

OpenStudy (nincompoop):

\[\lim_{x \rightarrow 0}\frac{ 2x^2-5x-3 }{ x-1 }\] can you factor the top part?

OpenStudy (anonymous):

So I would plug in -1.9 for all the x's?

OpenStudy (nincompoop):

you can if you want to see the value visually but we can do some algebraic manipulation to solve for it as well.

OpenStudy (anonymous):

I got -15.24444444444 doing it that way.

OpenStudy (nincompoop):

-0.9 -0.99 -0.999 -0.9999 NOT -1.99 also from the left side (negative side), you are going to use -1.01 -1.001 -1.0001 -1.00001

OpenStudy (anonymous):

I'm a bit confused. Sorry.

OpenStudy (anonymous):

So the 2x^2-5x-3

OpenStudy (anonymous):

(2x-3)(x+1)?

OpenStudy (anonymous):

I can't of anything else that would go into 3

OpenStudy (anonymous):

I see now.

OpenStudy (anonymous):

Do I need to set them to =0?

OpenStudy (anonymous):

I do not think so.

OpenStudy (anonymous):

Neither have (x+1)

OpenStudy (anonymous):

So nothing cancels out?

OpenStudy (anonymous):

3,-1

OpenStudy (anonymous):

X-3=0 x=3

OpenStudy (anonymous):

x+1=0 x=-1

OpenStudy (nincompoop):

erm

OpenStudy (anonymous):

so the slant would be x=-3?

OpenStudy (anonymous):

And there isn't a horizontal asymptote,right?

OpenStudy (nincompoop):

there might

OpenStudy (nincompoop):

your horizontal asymptote is -3, your vertical is -1 it is -1 because if you make x= -1 you will get a zero denominator

OpenStudy (anonymous):

Okay